OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)). -40x + 24 -40x + 31 -40x + 64 -40x + 70

OpenStudy (anonymous):

f(g(x))=-4(10x-6)+7=-40x+24+7=-40x+31 simply replace x in f(x) by the function g(x)...